# MaNGA DynPop – VII. A Unified Bulge-Disk-Halo Model for Explaining Diversity in Circular Velocity Curves of 6000 Spiral and Early-Type Galaxies

### v1

@@ 3.3 Bulge-disk-halo model to quantify the shape of CVCs
where $\Gamma(a)$ is the gamma function, $\Gamma(a,x)$ is the upper incomplete one, and $Q(a,x)$ is the regularized upper incomplete gamma function (Olver et al. 2010, equation 8.2.4). Various approximations have been proposed to calculate $b_{n}$ (e.g., Capaccioli, 1989; Prugniel & Simien, 1997; Ciotti & Bertin, 1999; Cappellari, 2023). However, one can also compute $b_{n}$ using the special function $Q^{-1}(a,s)$ giving the inverse of the regularized upper incomplete gamma function. This provides the solution for $z$ of $s=Q(a,z)$. From Equation (missing) 13 one can then just write
@@
The readers are referred to Vitral & Mamon (2020) for a summary of different approximations for the deprojection of Sersic profile. For convenience, we assume a total stellar mass $M_{\rm\ast}$ and a projected half-light radius $R_{\rm e}$, which will finally be removed in the normalized CVC. We use the bulge-to-total ratio $B/T$, $R_{\rm e}$, and the $R_{\rm e,bulge}/R_{\rm e,disk}$ ratio to calculate $R_{\rm e,bulge}$ and $R_{\rm e,disk}$. Then we can obtain $M_{\rm\infty,bulge}=M_{\rm\ast}\times{\rm B/T}$ and $M_{\rm\infty,disk}=M_{\rm\ast}\times{\rm(1-B/T)}$ assuming spatially constant stellar mass-to-light ratio, and further the $M_{\rm bulge}(r)$ and $M_{\rm disk}(r)$ profiles from Equation (missing) 17. Unlike previous studies that adopt different disk shapes depending on morphological types (e.g. a spherical disk for ETGs and an infinitely thin disk for LTGs, as discussed in section 2.1 of Dutton et al., 2011), we assume a spherical disk for all galaxies. This assumption avoids introducing discontinuities by making the disk shape changing from spherical to infinitely thin depend on morphologies. Although the spherical disk is a crude approximation, our toy model remains meaningful as it adopts various fitting functions of model parameters and does not pretend to be quantitatively accurate. Moreover, adopting an alternative extreme assumption that all galaxies have infinitely thin disk (Freeman, 1970) leads to a similar level of accuracy in our model (see Appendix A and Figure 7), which justifies that the choice of a universal disk shape assumption does not significantly affect the results. For the NFW dark matter halo with two free parameters ($r_{\rm s}$ and $\rho_{\rm s}$), we use the $r_{\rm s}/R_{\rm e}$ value to estimate $r_{\rm s}$ and use the dark matter fraction within a sphere of $R_{\rm e}$, defined as
@@
The $B/T-\lg\,\sigma_{\rm e}$, $\lg\,n_{\rm Ser,bulge}-\lg\,\sigma_{\rm e}$, and $f_{\rm DM}(<R_{\rm e})-\lg\,\sigma_{\rm e}$ (Lu et al., 2024, figure 2) relations show strong dependence on $\sigma_{\rm e}$, while $R_{\rm e,bulge}/R_{\rm e,disk}$ varies little with $\sigma_{\rm e}$. Although the median relation of $\lg\,(r_{\rm s}/R_{\rm e})-\lg\,\sigma_{\rm e}$ increases with higher $\sigma_{\rm e}$, however, the minimum $r_{\rm s}$ (for the median relation) is at least five times larger than $R_{\rm e}$, making the dark matter density profile within the kinematic data range is approximately power-law. Thus we use the median values $\lg\,(R_{\rm e,bulge}/R_{\rm e,disk})=-0.37$ and $\lg\,(r_{\rm s}/R_{\rm e})=0.74$ (the median value for $\lg\,(\sigma_{\rm e}/{\rm km\,s^{-1}})<2.1$, below which dark matter fractions become more significant) instead of the scaling relations for these two parameters.
@@
An interesting fact is that, given a velocity dispersion $\sigma_{\rm e}$, we can estimate the amplitude $V_{\rm circ}(R_{\rm e}^{\rm maj})$ (or $V_{\rm circ}^{\rm max}$) of CVC with an error of 9% (or 8%) and predict the CVC shape with nearly zero bias in most cases. This demonstrates that $\sigma_{\rm e}$ is a good proxy for the inner gravitational potential. Especially when IFU observations for high-redshift galaxies are quite expensive, one can apply the empirical aperture correction for velocity dispersion (Zhu et al., 2023a) to obtain $\sigma_{\rm e}$ as the proxy for inner gravitational potential.
@@
Figure 4: Scaling relations between structural parameters ($B/T$, $n_{\rm Ser,bulge}$, $R_{\rm e,bulge}/R_{\rm e,disk}$, $r_{\rm s}/R_{\rm e}$) and $\sigma_{\rm e}$ for the $\rm Qual\geqslant 1$ galaxies. $B/T$, $n_{\rm Ser,bulge}$, and $R_{\rm e,bulge}/R_{\rm e,disk}$ are taken from Domínguez Sánchez et al. (2022) (see Section 2.2 for definitions). $r_{\rm s}/R_{\rm e}$ is the ratio between scale radius of dark halo and $R_{\rm e}$ of the whole galaxy, where $r_{\rm s}$ is estimated from the Chabrier IMF-based stellar mass, the stellar mass-to-halo mass relation in Moster et al. (2013), and the mass-concentration relation in Dutton & Macciò (2014). The blue contours are the kernel density estimate for the galaxy distribution, while the black dashed curve and the gray shaded region represent the median value and [16th, 84th] percentile of values, respectively. The best-fitting relations to the median trends (Equation (missing) 21 and Equation (missing) 22) are shown in red solid curves. The horizontal red dashed line in the third panel represents the median value of $\lg\,(R_{\rm e,bulge}/R_{\rm e,disk})$ for the entire sample, while the horizontal red dashed line in the fourth panel represents the median value of $\lg\,(r_{\rm s}/R_{\rm e})$ for galaxies with $\lg\,(\sigma_{\rm e}/{\rm km\,s^{-1}})<2.1$.
@@
Figure 5: Comparisons between the BDH model-predicted shape of normalized CVCs (solid curves) and those derived from JAM (dashed curves) for four different velocity dispersion ($\sigma_{\rm e}$) bins, indicated by different color. In the left panel, we show the median profile of the JAM-derived CVCs for each $\sigma_{\rm e}$ bin, while the BDH model predictions are calculated at the bin centers using the scaling relations (three relations as a function of $\sigma_{\rm e}$, and two are constants) presented in the left panel. The ratio between the JAM-derived and BDH-predicted CVCs, along with the associated error (represented by the [16th, 84th] percentiles of the JAM-derived CVCs), is shown in the right panel.
@@
Figure 6: The distributions of $B/T$ (top left panel), $f_{\rm DM}(<R_{\rm e})$ (top right panel, taken from figure 16 in Zhu et al., 2024), $\lg\,n_{\rm Ser,bulge}$ (bottom left panel), and the median CVC within a given mass and size bin (bottom right panel) are shown on the $M_{\rm JAM}-R_{\rm e}^{\rm maj}$ plane. The distributions are smoothed by the loess software (Cappellari et al., 2013b) with frac=0.05. In all panels, the dashed lines correspond to effective velocity dispersion $\sigma_{\rm e}=$ 50, 100, 200, 300, 400, and 500 $\rm km\,s^{-1}$ from left to right, which are calculated using the scalar virial equation $\sigma_{\mathrm{e}}^{2}\equiv GM_{\rm JAM}/(5R_{\rm e}^{\rm maj})$ (with factor 5 from Cappellari et al. 2006). The magenta curve shows the zone of exclusion (ZOE) defined in (Cappellari et al., 2013b, equation 4), with the ZOE above $M_{\rm JAM}=2\times 10^{10}{\rm M_{\odot}}$ is approximately $R_{\rm e}^{\rm maj}\propto M_{\rm JAM}^{0.75}$. The gray contours show the kernel density estimate for the galaxy distribution. Bottom right panel: The inset panels in the bottom right panel show the median (red solid) and the [16th, 84th] percentiles (red shaded region), as well as the median FWHM of PSF (vertical shaded gray region). All the inset panels have the same physical scales as the bottom right panel, while the ticks and labels are shown in the upper left inset panel.
@@ 3.4 CVC shape on the mass-size plane
We revisit galaxy evolution (focusing on the CVC shapes) on the mass-size plane in Figure 6, in which we show the distributions of $B/T$, $f_{\rm DM}(<R_{\rm e})$, $n_{\rm Ser,bulge}$, and the median CVC for galaxies in each mass-size bin (corresponding to each inset panel). The bottom right panel reveals clear trends:
@@
As discussed in previous sections, the shape of CVCs is determined by $B/T$, $f_{\rm DM}(<R_{\rm e})$, and $n_{\rm Ser,bulge}$. Consequently, the evolution of CVC shapes can be understood through the evolution of these parameters. As suggested by the IFU results (Cappellari et al., 2013b; Cappellari, 2016; Zhu et al., 2024) and the observations of high-redshift ETGs (van der Wel et al., 2008; van Dokkum et al., 2015; Derkenne et al., 2021), galaxy evolution on the mass-size plane can be explained by the two-phase evolutionary scenario: (i) gas accretion-induced star formation and (ii) dry mergers.
When disk galaxies form (i.e., the formation of stellar disks; Mo et al., 1998), they typically exhibit small bulges (low $\sigma_{\rm e}$ and low $B/T$). These young disk galaxies tend to host pseudobulges (e.g. Hu et al., 2024), which typically exhibit exponential-like profiles with Sersic indices of $n_{\rm Ser,bulge}\approx 1-2$ (Kormendy & Kennicutt, 2004). Subsequent inner star formation triggered by gas accretion, drives bulge growth, resulting in a slightly higher Sersic index for the bulge and a lower central dark matter fraction. This transformation causes galaxies to shift their CVC shapes from rising to declining as their bulges and $\sigma_{\rm e}$ increase. The size dependence (quantified by the half-light radius) is primarily explained by the fact that, at fixed mass, galaxies with higher $B/T$ are smaller, and contain less dark matter within their smaller size (lower dark matter fraction). The trend differs for high $\sigma_{\rm e}$ ETGs that lie parallel to the ZOE, where dry mergers dominate mass and size growth. Along the direction of the ZOE, dry mergers mildly increase $B/T$ and have minimal impact on $f_{\rm DM}(<R_{\rm e})$. Although $n_{\rm Ser,bulge}$ changes more significantly, it plays a relatively less important role in determining CVC shape (see Section 3.3). As a result, these high $\sigma_{\rm e}$ galaxies exhibit nearly the same CVC shape.
@@ 4 Conclusions
In the seventh paper of the MaNGA DynPop project, we derive the CVCs (or intrinsic RCs) for 6000 nearby galaxies from stellar dynamical models in (Zhu et al., 2023b). The amplitude and shape of CVCs (corrected for the inclination angle and seeing effects) are closely related to the inner gravitational potential of galaxies, providing key insights into galaxy formation and evolution scenarios. By combining these with spatially resolved stellar population properties (Lu et al., 2023), we study the scaling relations between CVCs (shape and amplitude) and other galaxy properties. We propose a simple BDH model that includes a bulge, a disk and a dark matter halo to predict the shape of CVCs across different galaxy types.
The main conclusions are summarized as below.
@@
  • The amplitude of CVCs, characterized by the circular velocity at the half-light radius, $V_{\rm circ}(R_{\rm e}^{\rm maj})$, or the maximum circular velocity within the FoV, $V_{\rm circ}^{\rm max}$, is linearly related to the velocity dispersion within one effective radius $\sigma_{\rm e}$, linking the Tully-Fisher relation (Tully & Fisher, 1977) and the Faber-Jackson relation (Faber & Jackson, 1976). The best-fitting relations are $V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 1.65\times\sigma_{\rm e}$ and $V_{\rm circ}^{\rm max}\approx 1.75\times\sigma_{\rm e}$ with a small error of 9% or 8% (Figure 2), respectively.
@@
  • The shape of CVCs (rising, flat, and declining) correlates with galaxy dynamical and stellar population properties: galaxies with declining CVCs are massive, early-type, early-quenched, old, metal-rich, and exhibit high $B/T$, low spin, high velocity dispersion, and flat age gradient, while galaxies with rising CVCs exhibit the opposite properties (Figure 3).
@@
  • We propose a BDH model, which includes a Sersic bulge, an exponential disk, and a NFW dark matter halo, to quantify the shape of CVCs. In this model, given the bulge-to-total ratio $B/T$, the dark matter fraction within $f_{\rm DM}(<R_{\rm e})$, the Sersic index of the bulge component $n_{\rm Ser,bulge}$, the ratio between effective radii of bulge and disk $R_{\rm e,bulge}/R_{\rm e,disk}$, and the ratio between the scale radius of NFW halo and the luminous half-light radius $r_{\rm s}/R_{\rm e}$, one can predict the shape of CVCs. We test this model for galaxies within different $\sigma_{\rm e}$ bins, in which the free parameters are predicted from their empirical scaling relations with $\sigma_{\rm e}$ (see Figure 4 and figure 2 in Lu et al., 2024), finding a nearly unbiased consistency with JAM-derived CVC shapes (Figure 5). The BDH model quantitatively confirms that the shape of CVCs is mainly driven by $B/T$, $f_{\rm DM}(R_{\rm e)}$, and $n_{\rm Ser,bulge}$: galaxies with larger $B/T$, lower $f_{\rm DM}(R_{\rm e})$, and larger $n_{\rm Ser,bulge}$ tend to have declining CVCs, while rising CVCs are found in galaxies with opposite properties.
@@
  • The evolution of CVC shapes on the mass-size plane is closely related to the evolution of $B/T$, $f_{\rm DM}(R_{\rm e})$, and $n_{\rm Ser,bulge}$, supporting the scenario of two evolutionary channels (e.g. Cappellari et al., 2013b; van Dokkum et al., 2015; Zhu et al., 2024): (i) in situ star formation (through gas accretion or gas-rich minor mergers) moving galaxies from left to right (CVC shape from rising to declining), inducing bulge growth, reducing central dark matter fraction, and increasing the Sersic index of bulge; (ii) dry mergers moving gas-poor galaxies along the constant $\sigma_{\rm e}$ lines upwards, mildly increasing $B/T$ and the Sersic index of bulge, leaving the central dark matter fraction unchanged, and finally leading to the same CVC shapes along the direction of the ZOE (Figure 6).
@@
We will release all data derived from the stellar dynamical models, including 3D mass distributions (mass density profiles), 2D mass distributions (surface mass density maps), and circular velocity curves. One can use the MGE coefficients and best-fitting free parameters taken from the catalog (Zhu et al., 2023b) to derive the mass distributions of stellar, dark matter, and total components. The database and a Python script for the calculation will be provided on the website of MaNGA DynPop (https:manga-dynpop.github.io).

### v2

@@ 3.3 A Bulge-Disk-Halo model to quantify the shape of CVCs
where $\Gamma(a)$ is the gamma function, $\Gamma(a,x)$ is the upper incomplete one, and $Q(a,x)$ is the regularized upper incomplete gamma function (Olver et al. 2010, equation 8.2.4). Various approximations have been proposed to calculate $b_{n}$ (e.g., Capaccioli, 1989; Prugniel & Simien, 1997; Ciotti & Bertin, 1999; Cappellari, 2023). However, one can also compute $b_{n}$ using the special function $Q^{-1}(a,s)$ giving the inverse of the regularized upper incomplete gamma function. This provides the solution for $z$ of $s=Q(a,z)$. From Equation () 13 one can then just write
@@
Readers are referred to Vitral & Mamon (2020) for a summary of different approximations for the deprojection of Sersic profile. For convenience, we assume a total stellar mass $M_{\rm\ast}$ and a projected half-light radius $R_{\rm e}$, which will finally be removed in the normalized CVC. We use the bulge-to-total ratio $B/T$, $R_{\rm e}$, and the $R_{\rm e,bulge}/R_{\rm e,disk}$ ratio to calculate $R_{\rm e,bulge}$ and $R_{\rm e,disk}$. Then we can obtain $M_{\rm\infty,bulge}=M_{\rm\ast}\times{\rm B/T}$ and $M_{\rm\infty,disk}=M_{\rm\ast}\times{\rm(1-B/T)}$ assuming a spatially constant stellar mass-to-light ratio, and further the $M_{\rm bulge}(r)$ and $M_{\rm disk}(r)$ profiles from Equation () 17. Unlike previous studies that adopt different disk shapes depending on morphological types (e.g. a spherical disk for ETGs and an infinitely thin disk for LTGs, as discussed in section 2.1 of Dutton et al., 2011), we assume a spherical disk for all galaxies. This assumption avoids introducing discontinuities by making the disk shape changing from spherical to infinitely thin dependent on morphologies. Although the spherical disk is a crude approximation, our toy model remains meaningful as it adopts various fitting functions of model parameters and does not pretend to be quantitatively accurate. Moreover, adopting an alternative extreme assumption that all galaxies have an infinitely thin disk (Freeman, 1970) leads to a similar level of accuracy in our model (see Appendix A and Figure 7), which justifies that the choice of a universal disk-shape assumption does not significantly affect the results. For the NFW DM halo with two free parameters ($r_{\rm s}$ and $\rho_{\rm s}$), we use the $r_{\rm s}/R_{\rm e}$ value to estimate $r_{\rm s}$ and use the DM fraction within a sphere of $R_{\rm e}$, defined as
@@
The $B/T-\lg\,\sigma_{\rm e}$, $\lg\,n_{\rm Ser,bulge}-\lg\,\sigma_{\rm e}$, and $f_{\rm DM}(<R_{\rm e})-\lg\,\sigma_{\rm e}$ (Lu et al., 2024, figure 2) relations show strong dependence on $\sigma_{\rm e}$, while $R_{\rm e,bulge}/R_{\rm e,disk}$ varies little with $\sigma_{\rm e}$. Although the median relation of $\lg\,(r_{\rm s}/R_{\rm e})-\lg\,\sigma_{\rm e}$ increases with higher $\sigma_{\rm e}$, however, the minimum $r_{\rm s}$ (for the median relation) is at least 5 times larger than $R_{\rm e}$, meaning the DM density profile within the kinematic data range is approximately power law. Thus we use the median values $\lg\,(R_{\rm e,bulge}/R_{\rm e,disk})=-0.37$ and $\lg\,(r_{\rm s}/R_{\rm e})=0.74$ (the median value for $\lg\,(\sigma_{\rm e}/{\rm km\,s^{-1}})<2.1$, below which DM fractions become more significant) instead of the scaling relations for these two parameters.
@@
An interesting fact is that, given a velocity dispersion $\sigma_{\rm e}$, we can estimate the amplitude $V_{\rm circ}(R_{\rm e}^{\rm maj})$ (or $V_{\rm circ}^{\rm max}$) of CVC with an error of 9% (or 8%) and predict the CVC shape with nearly zero bias in most cases. This demonstrates that $\sigma_{\rm e}$ is a good proxy for the inner gravitational potential. In particular, when IFU observations for high-redshift galaxies are quite expensive, one can apply the empirical aperture correction for velocity dispersion (Zhu et al., 2023a) to obtain $\sigma_{\rm e}$ as the proxy for inner gravitational potential.
@@
Figure 4: Scaling relations between structural parameters ($B/T$, $n_{\rm Ser,bulge}$, $R_{\rm e,bulge}/R_{\rm e,disk}$, $r_{\rm s}/R_{\rm e}$) and $\sigma_{\rm e}$ for the $\rm Qual\geqslant 1$ galaxies. $B/T$, $n_{\rm Ser,bulge}$, and $R_{\rm e,bulge}/R_{\rm e,disk}$ are taken from Domínguez Sánchez et al. (2022) (see Section 2.2 for definitions). $r_{\rm s}/R_{\rm e}$ is the ratio between the dark halo scale radius and $R_{\rm e}$ of the whole galaxy, where $r_{\rm s}$ is estimated from the Chabrier IMF-based stellar mass, the stellar mass-to-halo mass relation in Moster et al. (2013), and the mass-concentration relation in Dutton & Macciò (2014). The blue contours are the kernel density estimate for the galaxy distribution, while the black dashed curve and the gray shaded region represent the median value and [16th, 84th] percentile of values, respectively. The best-fitting relations to the median trends (Equation () 21 and Equation () 22) are shown in red solid curves. The horizontal red dashed line in the third panel represents the median value of $\lg\,(R_{\rm e,bulge}/R_{\rm e,disk})$ for the entire sample, while the horizontal red dashed line in the fourth panel represents the median value of $\lg\,(r_{\rm s}/R_{\rm e})$ for galaxies with $\lg\,(\sigma_{\rm e}/{\rm km\,s^{-1}})<2.1$.
@@
Figure 5: Comparisons between the BDH model-predicted shape of normalized CVCs (solid curves) and those derived from JAM (dashed curves) for four different velocity dispersion ($\sigma_{\rm e}$) bins, indicated by different colors. In the left panel, we show the median profile of the JAM-derived CVCs for each $\sigma_{\rm e}$ bin, while the BDH model predictions are calculated at the bin centers using the scaling relations (three relations as a function of $\sigma_{\rm e}$, and two are constants) presented in the left panel. The ratio between the JAM-derived and BDH-predicted CVCs, along with the associated error (represented by the [16th, 84th] percentiles of the JAM-derived CVCs), is shown in the right panel.
@@
Figure 6: The distributions of $B/T$ (top-left panel), $\lg\,n_{\rm Ser,bulge}$ (top-right panel), $T_{50}/T_{90}$ (bottom-left panel), and the median CVC within a given mass and size bin (bottom-right panel) are shown on the $M_{\rm JAM}-R_{\rm e}^{\rm maj}$ plane. Distributions of $f_{\rm DM}(<R_{\rm e})$ can be found in Zhu et al. (2024, top-right panel of figure 16). The distributions are smoothed by the loess software (Cappellari et al., 2013b) with frac=0.05. In all panels, the dashed lines correspond to effective velocity dispersion $\sigma_{\rm e}=$ 50, 100, 200, 300, 400, and 500 $\rm km\,s^{-1}$ from left to right, which are calculated using the scalar virial equation $\sigma_{\mathrm{e}}^{2}\equiv GM_{\rm JAM}/(5R_{\rm e}^{\rm maj})$ (with factor 5 from Cappellari et al. 2006). The magenta curve shows the zone of exclusion (ZOE) defined in (Cappellari et al., 2013b, equation 4), with the ZOE above $M_{\rm JAM}=2\times 10^{10}{\rm M_{\odot}}$ is approximately $R_{\rm e}^{\rm maj}\propto M_{\rm JAM}^{0.75}$. The gray contours show the kernel density estimate for the galaxy distribution. Bottom-right panel: The inset panels in the bottom-right panel show the median (red solid) and the [16th, 84th] percentiles (red shaded region), as well as the median FWHM of PSF (vertical shaded gray region). All the inset panels have the same physical scales as the bottom-right panel, while the ticks and labels are shown in the upper-left inset panel.
@@ 3.4 CVC shape on the mass-size plane
It has been suggested by the IFU results of nearby ETGs (Cappellari et al., 2013b; Cappellari, 2016; Zhu et al., 2024) and the observations of high-redshift ETGs (van der Wel et al., 2008; van Dokkum et al., 2015; Derkenne et al., 2021) that galaxy evolution on the mass-size plane follows a simple two-phase scenario (Oser et al., 2010; Tonini et al., 2016): (i) in situ star formation and (ii) ex situ accretion by dry mergers (minor or major). In situ star formation, whether triggered by infalling cold gas (e.g. cold streams or gas-rich minor mergers) or pure secular evolution without gas replenishment, mildly increases galaxy sizes when building up stellar masses (Oser et al., 2010). In contrast, dry major mergers lead to proportional increases in both galaxy sizes and masses (while maintaining nearly constant $\sigma_{e}$), whereas dry minor mergers result in more significant size growth, typically by a factor of 4 for mass doubling (Naab et al., 2009).
The two-phase scenario is not intended as a full picture of galaxy formation but instead serves as a framework for understanding the physical processes in a straightforward way. As discussed in previous sections, the shape of CVCs is determined by $B/T$, $f_{\rm DM}(<R_{\rm e})$, and $n_{\rm Ser,bulge}$. Consequently, the evolution of CVC shapes on the mass-size plane is related to the evolution of these parameters. Since the distribution of $f_{\rm DM}(<R_{\rm e})$ has been presented in Zhu et al. (2024, figure 16), we show the distributions of $B/T$, $n_{\rm Ser,bulge}$, $T_{50}/T_{90}$, and the median CVC for galaxies in each mass-size bin (corresponding to each inset panel) in Figure 6. $T_{50}/T_{90}$ is the ratio between cosmic times when 50% and 90% stellar masses are formed. The bottom-right panel reveals clear trends of CVC shapes:
@@
When disk galaxies form (i.e., the formation of stellar disks; Mo et al., 1998), they typically exhibit small bulges (low $\sigma_{\rm e}$ and low $B/T$). These young disk galaxies tend to host pseudobulges (e.g. Hu et al., 2024), which are induced by secular evolution and typically exhibit exponential-like profiles with Sersic indices of $n_{\rm Ser,bulge}\approx 1-2$ (Kormendy & Kennicutt, 2004). Inner in situ star formation, whether driven by secular evolution or enhanced through gas accretion, drives bulge growth, resulting in a slightly higher Sersic index for the bulge and a lower central DM fraction. This transformation causes galaxies to shift their CVC shapes from rising to declining as their bulges and $\sigma_{\rm e}$ increase. The size dependence (quantified by the half-light radius) is primarily explained by the fact that, at fixed mass, galaxies with higher $B/T$ are smaller, and contain less DM within their smaller size (lower DM fraction). The trend differs for high-$\sigma_{\rm e}$ ETGs that lie parallel to the ZOE, where dry mergers dominate mass and size growth. Along the direction of the ZOE, dry mergers mildly increase $B/T$ and have minimal impact on $f_{\rm DM}(<R_{\rm e})$. Although $n_{\rm Ser,bulge}$ changes more significantly, it plays a relatively less important role in determining CVC shape (see Section 3.3). As a result, these high-$\sigma_{\rm e}$ galaxies exhibit nearly the same CVC shape.
In the bottom-left panel of Figure 6, the star-forming galaxies at $10^{10}-10^{11}{\rm M_{\odot}}$ exhibit relatively smaller $T_{50}/T_{90}$ values compared to other star-forming galaxies with lower or higher masses. This is likely an in situ star formation region dominated by pure secular evolution. Among the in situ star formation processes, secular evolution dominates at intermediate masses ($10^{10}-10^{11}{\rm M_{\odot}}$), as gas accretion rates/merger rates are typically higher at lower/higher masses (Hopkins et al., 2010; Tonini et al., 2016). In the absence of gas replenishment, secular evolution tends to form stars at progressively slower rates (due to declining gas densities), resulting in the observed smaller $T_{50}/T_{90}$ ratios.
Other mechanisms are also potentially responsible for the evolution of galaxy mass and size, such as tidal heating and tidal stripping. Tidal heating puffs up the galaxies to form ultra-diffuse galaxies (Jones et al., 2021; Fielder et al., 2024), while tidal stripping leads to ultra compact dwarfs (Penny et al., 2014; Mayes et al., 2021). However, these environmental effects only work for low-mass satellite galaxies in a dense environment, which is not suitable for our sample dominated by central galaxies (Zhu et al., 2024, section 2.5).
@@ 4 Conclusions
In the seventh paper of the MaNGA DynPop project, we derive the CVCs (or intrinsic RCs) for 6000 nearby galaxies from stellar dynamical models in (Zhu et al., 2023b). The amplitude and shape of CVCs (corrected for the inclination angle and seeing effects) are closely related to the inner gravitational potential of galaxies, providing key insights into galaxy formation and evolution scenarios. By combining these with spatially resolved stellar population properties (Lu et al., 2023), we study the scaling relations between CVCs (shape and amplitude) and other galaxy properties. We propose a simple BDH model that includes a bulge, a disk, and a DM halo to predict the shape of CVCs across different galaxy types.
The main conclusions are summarized below.
@@
  • The amplitude of CVCs, characterized by the circular velocity at the half-light radius, $V_{\rm circ}(R_{\rm e}^{\rm maj})$, or the maximum circular velocity within the FoV, $V_{\rm circ}^{\rm max}$, is linearly related to the velocity dispersion within one effective radius $\sigma_{\rm e}$, linking the Tully-Fisher relation (Tully & Fisher, 1977) and the Faber-Jackson relation (Faber & Jackson, 1976). The best-fitting relations are $V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 1.62\times\sigma_{\rm e}$ and $V_{\rm circ}^{\rm max}\approx 1.72\times\sigma_{\rm e}$ with a small error of 7% (Figure 2), respectively.
@@
  • The shape of CVCs (rising, flat, and declining) correlates with galaxy dynamical and stellar population properties: galaxies with declining CVCs are massive, early-type, early-quenched, old, metal-rich, and exhibit high $B/T$, low spin, high velocity dispersion, and a flat age gradient, while galaxies with rising CVCs exhibit the opposite properties (Figure 3).
@@
  • We propose a BDH model, which includes a Sersic bulge, an exponential disk, and a NFW DM halo, to quantify the shape of CVCs. In this model, given the bulge-to-total ratio $B/T$, the dark matter fraction within $f_{\rm DM}(<R_{\rm e})$, the Sersic index of the bulge component $n_{\rm Ser,bulge}$, the ratio between the bulge and disk effective radii $R_{\rm e,bulge}/R_{\rm e,disk}$, and the ratio between the scale radius of NFW halo and the luminous half-light radius $r_{\rm s}/R_{\rm e}$, one can predict the shape of CVCs. We test this model for galaxies within different $\sigma_{\rm e}$ bins, in which the free parameters are predicted from their empirical scaling relations with $\sigma_{\rm e}$ (see Figure 4 and figure 2 in Lu et al., 2024), finding a nearly unbiased consistency with JAM-derived CVC shapes (Figure 5). The BDH model quantitatively confirms that the shape of CVCs is mainly driven by $B/T$, $f_{\rm DM}(R_{\rm e)}$, and $n_{\rm Ser,bulge}$: galaxies with a larger $B/T$, lower $f_{\rm DM}(R_{\rm e})$, and larger $n_{\rm Ser,bulge}$ tend to have declining CVCs, while rising CVCs are found in galaxies with the opposite properties.
@@
  • The evolution of CVC shapes on the mass-size plane is closely related to the evolution of $B/T$, $f_{\rm DM}(R_{\rm e})$, and $n_{\rm Ser,bulge}$, supporting the scenario of two evolutionary channels (e.g. Cappellari et al., 2013b; van Dokkum et al., 2015; Zhu et al., 2024): (i) in situ star formation (through gas accretion, gas-rich minor mergers, or secular evolution) moving galaxies from left to right (CVC shape from rising to declining), inducing bulge growth, reducing the central DM fraction, and increasing the Sersic index of bulges; (ii) dry mergers moving gas-poor galaxies along the constant $\sigma_{\rm e}$ lines upwards, mildly increasing $B/T$ and the Sersic index of bulges, leaving the central DM fraction unchanged, and finally leading to the same CVC shapes along the direction of the ZOE (Figure 6).
@@
We will release all data derived from the stellar dynamical models, including 3D mass distributions (mass density profiles), 2D mass distributions (surface mass density maps), and circular velocity curves. One can use the MGE coefficients and best-fitting free parameters taken from the catalog (Zhu et al., 2023b) to derive the mass distributions of stellar, DM, and total components. The database and a Python script for the calculation will be provided on the website of MaNGA DynPop (https:manga-dynpop.github.io).